# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

### v1

@@ V-C Results of Evasion Attacks
Moreover, IF-based methods generally exhibit moderate robustness, with DUA demonstrating relatively higher robustness. DUA enhances auditing effectiveness by injecting additional noise into the dataset. This mechanism improves robustness but also makes the audit results more susceptible to forgery. We will further discuss the details in Section V-E.
TABLE VI: Results of forgery attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@
IF-based MIA Score ($\downarrow$) 0.54 0.66 0.72 0.54 0.71 0.71
Rapid Score ($\downarrow$) 0.22 0.79 0.89 0.30 0.86 0.88
DI p-value ($\uparrow$) 0.01 $10^{-13}$ $10^{-5}$ $10^{-23}$ 0.97 0.01
Diff ($\downarrow$) 0.79 1.35 0.98 1.61 -0.81 0.09
DUA Cost (logits) (%, $\uparrow$) 100.00 0.13 0.13 0.18 0.13 0.24
Cost (label) (%, $\uparrow$) 100.00 0.13 0.13 0.24 0.13 0.24
EF-based DVBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 10.01 29.92 98.37 65.60 69.77 96.77
UBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 11.73 86.55 100.00 71.08 95.69 100.00
ZeroMark p-value ($\uparrow$) 1.00 1.00 1.00 0.99 0.01 1.00
WSR (%, $\downarrow$) 10.01 29.93 98.44 68.28 71.68 96.88
DW WSR (%, $\downarrow$) 11.73 67.00 89.40 66.37 73.72 89.36
V-D Results of Hybrid Evasion Attacks
We also explore combining evasion attack methods from different phases to enhance overall attack effectiveness. Specifically, we hereby evaluate three types of hybrid evasion attacks: (1) HA1: AutoEncoder, Adversarial Training, and Reprogramming. (2) HA2: WaveletFilter, DP-SGD, and Randomized Smoothing. (3) HA3: WaveletFilter and Reprogramming. The results are shown in Table V.
Takeaway 4: Combining attacks from different phases can significantly enhance the attack effectiveness.
@@ V-D Results of Hybrid Evasion Attacks
V-E Results of Forgery Attacks
For forgery attacks, we assume that the adversary wants to make their own dataset ($i.e.$, the attack auxiliary dataset) to be falsely identified as the ‘trained dataset’ of other independent models. We use the model trained on the unmarked release dataset as the independent model. The results of white-box forgery attacks are in Table VI, and the experiments in the black-box scenario can be found in Appendix -G.
Takeaway 5: Existing dataset auditing methods are overall vulnerable to forgery attacks.
Takeaway 6: Dataset auditing methods that involve more complex decision-making and auditing processes are more resistant to forgery.
V-F Attack Overhead
V-G Summary of Evaluated Dataset Auditing Methods
To provide a holistic comparison of the evaluated dataset auditing methods, we summarize their relative performance across five key dimensions: Robustness, Distinctiveness, Fidelity, Stealthiness, and Efficiency according to the design objectives in Section III-C. Figure 4 represents this comparison using a radar chart, where scores closer to the outer edge (1.0) indicate better performance in that dimension.
VI Discussion
VI-A Related Works
Since dataset auditing in DL is an emerging topic, there is only a small number of works surveying this domain. Asswad et al. [2] highlighted the importance of protecting data ownership and reviewed several SOTA techniques in a wide range of data ownership. Chandrasekaran et al. [9] provided a comprehensive study about machine learning governance, including data ownership in machine learning. Recently, Du et al. [18] presented the first systematic SoK regarding dataset auditing, primarily focusing on surveying existing solutions. From the broader perspective of AI copyright protection, the majority of existing work focuses on watermarking techniques for models [68], generated content [69, 88], and traditional multimedia [54]. Lukas et al. [57] present a toolbox to evaluate the robustness of DNN watermarking, which is fundamentally different from dataset auditing.
VI-B Dataset Auditing for Generative Models
Given these challenges, existing auditing methods are not maturely applicable and may yield unstable or misleading results when evaluated under adversarial settings. Therefore, we caution that rushing into adversarial robustness evaluations in such an ill-defined setting without reliable baselines may misrepresent the true difficulty of the problem or misguide the community toward premature conclusions. We leave this as an important direction for future work.
VI-C Potential Future Direction
Future Direction for Dataset Auditing. The demonstrated vulnerability of existing dataset auditing methods underscores an urgent need for fundamentally more robust and distinctive techniques. Future work may focus on developing novel auditing mechanisms resilient to these attacks. (1) For evasion, this could involve exploring IFs that are harder to decouple from the model utility. For EF-based methods, future works may seek EFs that are deeply integrated into the model’s core functionality rather than easily removable or detectable trigger patterns. A certified robust mechanism is also urgently needed. (2) For forgery, our empirical study suggests that complicating the auditing process may lead to better resistance. Consequently, exploring more complex methods, such as multi-bit watermarks or cryptography [25, 75], may be a promising future direction to improve dataset auditing.
Limitations and Future Directions for Evaluating Dataset Auditing. While DATABench provides a comprehensive benchmark, there are still some limitations. First, this paper primarily focuses on dataset auditing regarding image datasets since they are currently the most widely studied and feasible methods. There have been some preliminary works studying auditing datasets of other modalities, such as text [41, 84], point clouds [83], and audio [51]. Providing a comprehensive evaluation of these works may be an important future work. Second, this paper does not cover all possible scenarios of dataset usage in DL. We primarily evaluate the two most important and widely adopted settings: training from scratch (Section V) and fine-tuning (Appendix -F). Third, DATABench does not thoroughly explore finer-grained threat models, such as adversaries with varying levels of capabilities (an adversary with very limited computational resources or data). An exhaustive study of the attacker’s capabilities would be a meaningful direction for future work.
VII Conclusion
This paper presented a comprehensive adversarial evaluation of dataset auditing in DL systems, addressing the critical gap between existing auditing methods and their resilience against dedicated attacks. We introduced a novel internal feature and external feature taxonomy to classify auditing techniques, formally defined evasion and forgery attack objectives, and proposed systematic attack strategies to realize them. These contributions led to the development of DATABench, a benchmark featuring 17 evasion and 5 forgery attacks. DATABench also provides an extensible toolbox to design new dataset auditing methods and attacks. This toolbox can also facilitate the unified evaluation of auditing methods. Our empirical experiments of 9 representative auditing methods against DATABench revealed some critical findings, $e.g.$, none of them demonstrated sufficient robustness or distinctiveness under adversarial settings. These results highlight the significant vulnerability of current dataset auditing methods and underscore the urgent need for developing novel and adversarially resilient dataset auditing methods.
Ethic Considerations

### v2

@@ Appendix E Discussion of the Model Utility Degradation
Table 7: Results of representative evasion attacks against transformer-based models trained on ImageNet. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack AutoEncoder AdvTraining-FGSM AdvTraining-Hybrid RS Reprogramming
IF-based MIA Test Acc. ($\%,\uparrow$) 74.88 72.76 64.54 65.56 64.70 69.84
Score ($\uparrow$) 0.86 0.86 0.72 0.72 0.00 0.00
Rapid Test Acc. ($\%,\uparrow$) 74.88 72.76 64.54 65.56 61.58 69.54
Score ($\uparrow$) 0.98 0.98 0.88 0.91 0.00 0.00
DI Test Acc. ($\%,\uparrow$) 74.88 72.76 64.54 65.56 63.20 68.26
p-value ($\downarrow$) $10^{-4}$ $10^{-12}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ 0.07
Diff ($\uparrow$) 0.62 0.61 0.50 0.69 0.50 0.31
DUA Test Acc. ($\%,\uparrow$) 75.74 73.18 64.36 66.70 70.52 69.70
Cost (logits) ($\%,\downarrow$) 30.60 69.23 100.00 99.80 47.14 35.63
Cost (label) ($\%,\downarrow$) 60.46 88.23 100.00 100.00 83.63 86.20
EF-based DVBW Test Acc. ($\%,\uparrow$) 73.94 70.28 62.82 65.08 60.16 67.66
p-value ($\downarrow$) 0.00 0.00 0.89 0.00 1.00 1.00
WSR ($\%,\uparrow$) 100.00 91.30 37.69 100.00 100.00 0.91
UBW-P Test Acc. ($\%,\uparrow$) 73.62 71.12 62.04 64.40 62.80 67.80
p-value ($\downarrow$) 0.00 0.00 0.00 0.00 1.00 1.00
WSR ($\%,\uparrow$) 98.34 51.64 98.28 98.44 97.86 34.81
UBW-C Test Acc. ($\%,\uparrow$) 73.36 71.54 64.50 66.38 64.22 66.34
p-value ($\downarrow$) 1.00 1.00 1.00 1.00 1.00 1.00
WSR ($\%,\uparrow$) 20.00 13.92 16.46 11.39 36.70 17.72
ZeroMark Test Acc. ($\%,\uparrow$) 71.50 69.00 59.00 60.98 51.50 66.60
p-value ($\downarrow$) 0.01 0.01 0.79 0.03 $10^{-5}$ 1.00
WSR ($\%,\uparrow$) 97.6 87.69 73.73 100.00 100.00 93.51
DW Test Acc. ($\%,\uparrow$) 73.94 72.34 64.00 65.46 70.40 67.62
WSR ($\%,\uparrow$) 0.00 18.00 0.00 1.99 0.00 0.00
Table 8: Results of representative evasion attacks in the fine-tuning scenario. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack AutoEncoder AdvTraining-Hybrid DP-SGD ($\epsilon=64$) RS Reprogramming
IF-based MIA Test Acc. ($\%,\uparrow$) 88.26 85.70 85.60 72.87 85.70 84.96
Score ($\uparrow$) 0.62 0.59 0.56 0.52 0.00 0.00
Rapid Test Acc. ($\%,\uparrow$) 88.26 85.70 85.60 72.87 85.70 86.24
Score ($\uparrow$) 0.62 0.65 0.68 0.48 0.33 0.27
DI Test Acc. ($\%,\uparrow$) 88.26 85.70 85.60 72.87 85.70 85.71
p-value ($\downarrow$) 0.06 0.55 0.10 0.05 $10^{-4}$ 0.11
Diff ($\uparrow$) 0.27 0.00 0.00 0.10 1.14 0.00
DUA Test Acc. ($\%,\uparrow$) 87.87 85.44 85.33 72.25 86.90 77.61
Cost (logits) ($\%,\downarrow$) 23.77 21.97 89.97 99.06 34.46 100.00
Cost (label) ($\%,\downarrow$) 79.03 84.60 99.91 89.09 89.71 100.00
EF-based DVBW Test Acc. ($\%,\uparrow$) 87.50 83.83 83.89 67.84 63.45 75.90
p-value ($\downarrow$) 0.00 1.00 0.00 0.00 1.00 1.0
WSR ($\%,\uparrow$) 100.00 20.01 99.92 100.00 100.00 10.74
UBW-P Test Acc. ($\%,\uparrow$) 87.66 84.39 83.10 68.07 80.06 85.71
p-value ($\downarrow$) 0.00 1.00 0.00 0.00 1.00 1.00
WSR ($\%,\uparrow$) 86.93 17.37 82.92 84.69 87.80 15.20
UBW-C Test Acc. ($\%,\uparrow$) 86.74 84.90 84.68 68.75 78.70 84.56
p-value ($\downarrow$) $10^{-143}$ $10^{-148}$ $^{-90} 1.00 1$.00 1.00
WSR ($\%,\uparrow$) 76.24 77.98 65.04 25.49 98.63 76.99
ZeroMark Test Acc. ($\%,\uparrow$) 87.81 83.39 83.19 71.45 65.26 85.56
p-value ($\downarrow$) $^{-98} 1.00 1$.00 1.00 1.00 1.00
WSR ($\%,\uparrow$) 100.00 15.72 99.64 100.00 100.00 10.68
DW Test Acc. ($\%,\uparrow$) 87.74 85.52 84.83 67.90 86.99 73.49
WSR ($\%,\uparrow$) 68.60 16.50 0.50 0.50 69.70 70.00
Table 9: Results of forgery attacks in the black-box scenario. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@
IF-based MIA Score ($\downarrow$) 0.54 0.54 0.55 0.53 0.55 0.55
Rapid Score ($\downarrow$) 0.22 0.16 0.28 0.15 0.25 0.13
DI p-value ($\uparrow$) 0.01 $10^{-48}$ $10^{-72}$ $10^{-14}$ 0.88 $10^{-4}$
Diff ($\downarrow$) 0.79 1.86 1.97 1.53 -0.42 0.41
DUA Cost (logits) (%, $\uparrow$) 100.00 0.13 0.37 0.41 0.19 0.13
Cost (label) (%, $\uparrow$) 100.00 1.27 1.71 0.69 0.19 0.13
EF-based DVBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 $10^{-144}$ 0.00
WSR (%, $\downarrow$) 10.01 16.87 18.53 35.12 23.48 33.25
UBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 $10^{-236}$ 0.00
WSR (%, $\downarrow$) 11.73 21.73 22.29 35.99 29.44 37.95
ZeroMark p-value ($\uparrow$) 1.00 0.99 1.00 0.99 1.00 1.00
WSR (%, $\downarrow$) 10.01 10.36 10.73 10.81 11.49 12.02
DW WSR (%, $\downarrow$) 11.73 20.68 19.49 35.15 25.28 34.35
Appendix F Experiments of Hybrid Evasion Attacks
We also explore combining evasion attack methods from different phases to enhance overall attack effectiveness. Specifically, we hereby evaluate three types of hybrid evasion attacks: (1) HA1: AutoEncoder, Adversarial Training, and Reprogramming. (2) HA2: WaveletFilter, DP-SGD, and Randomized Smoothing. (3) HA3: WaveletFilter and Reprogramming. The results are shown in Table 6.
Takeaway 6: Combining attacks from different phases can significantly enhance the attack effectiveness.
@@ Appendix F Experiments of Hybrid Evasion Attacks
Appendix G Experiments on Extra Models and Datasets
Experimental Settings. We conduct evaluations on the (100-class subset of) ImageNet dataset [17] using a different adversary’s model architecture, MobileViT [62], trained from scratch. We maintain the realistic assumption that if an auditor needed to train auxiliary models for their auditing method ($e.g.$, shadow models for MIA), they would use a standard architecture (ResNet-18), thus creating an architectural mismatch with the ViT suspect model.
Results. Consistent with our primary findings, the results in Table 7 show that existing dataset auditing methods remain highly vulnerable to evasion attacks in this more challenging setting. Both IF-based and EF-based auditing techniques frequently failed to correctly identify the usage of the ImageNet victim dataset when attacks like AutoEncoder-based denoising, adversarial training, or Reprogramming were employed. The architectural mismatch (transformer-based model and auditor’s CNN) and the increased dataset complexity do not confer robustness to the evaluated auditing methods.
Appendix H Evaluating Dataset Auditing in the Fine-tuning Scenario
Experimental Settings. Beyond training models from scratch, fine-tuning pre-trained models on a task-specific dataset is a common and efficient paradigm in modern DL. We simulate the fine-tuning scenario using a ResNet-18 pre-trained on ImageNet-1000 and then fine-tune it using CIFAR-10 as the target (victim) dataset. Other experimental settings remain consistent with our main experiments in Section 4.
Results. The outcomes of applying evasion attacks in the fine-tuning scenario are presented in Table 8. The results indicate that dataset auditing methods exhibit vulnerabilities similar to those observed when training from scratch. Both IF-based methods and EF-based methods frequently fail when confronted with evasion attacks such as DP-SGD, RS, or Reprogramming. This suggests that the inherent weaknesses of current auditing techniques are not limited to models trained from scratch but also extend to the fine-tuning paradigm.
Appendix I Forgery Attacks in the Black-box Scenario
This section evaluates forgery attacks when the adversary only possesses black-box access to the target model. Lacking parameter access hinders gradient-based attacks. Thus, the adversary may distill a substitute model from the target model using an auxiliary dataset. Then they can conduct forgery attacks on this substitute model, relying on the transferability of adversarial attacks to affect the original target model.
Experimental Settings. We simulate the black-box setting where the target model is the ResNet-18 and the substitute model employs the MobileNetV2 architecture[72]. All other settings remain consistent with the main experiments.
Results. From Table 9, the success rate of forgery attacks generally decreases in the black-box setting compared to the white-box scenario. However, despite the reduction, most dataset auditing methods ($e.g.$, DUA, DVBW, and UBW) remain susceptible to forgery even under these stricter black-box constraints. This indicates that even without direct model access, achieving dataset forgery remains a practical threat against a significant portion of current auditing techniques.
Appendix J Comparison with Other Related Data Protection Techniques
Traditional Data Watermarking. Traditional data watermarking techniques help track ownership and usage of data by embedding hidden markers ($i.e.$, watermark) into the data [44, 87]. The data owner can extract the watermark from the data and prove their ownership. Multiple solutions to embedding a watermark into data have been proposed, such as embedding watermarks into the frequency domain or utilizing DL models [9, 4]. Recently, some works have also explored using data watermarking to mark AI-generated content and models [86, 75, 45, 70]. However, traditional data watermarking techniques may not be feasible for dataset auditing since the watermark in the data can hardly be inherited by the DL models, and their owners will not disclose their training details (especially their training data).